OpenStudy (anonymous):

How do I solve 68% of 250? I need the steps, not the answer! I will give medal!

OpenStudy (anonymous):

68/100=x/250 then find x which is 68% of 250.

OpenStudy (anonymous):

@beloab I still do not understand

OpenStudy (anonymous):

% is a 100, so the full parts are 100 and 250. so x=68*250/100